# Language Inclusion for Boundedly-Ambiguous Vector Addition Systems is Decidable

### v3

@@ 3.2. Upper bounds
For a $k$-deterministic $d$-VASS one can build in exponential time a $(k\cdot d)$-dimensional downward-VASS which recognises the complement of its language.
@@
Before starting the proof, let us remark that it would seem natural to first build a deterministic $(k\cdot d)$-VASS equivalent to the input $k$-deterministic $d$-VASS and then apply construction from the proof of Theorem 15 to recognise its complement. However, it is not clear how to construct a deterministic $(k\cdot d)$-VASS equivalent to $k$-deterministic $d$-VASS, thus we compute directly a VASS recognising the complement of the input VASS language.
Let $V=(\Sigma,Q,T,c_{I},F)$ be a $k$-deterministic $d$-VASS. We aim to construct a $(k\cdot d)$-dimensional downward-VASS $V^{\prime}=(\Sigma,Q^{\prime},T^{\prime},c^{\prime}_{I},F^{\prime})$ such that $L(V^{\prime})=\Sigma^{*}\setminus L(V)$. In this proof, also, we strongly rely on the ideas introduced in the proof of Theorem 15. Recall that the idea of the construction is that $V^{\prime}$ simulates $k$ copies of $V$ which take care of different maximal runs of $V$. Then the accepting condition $F^{\prime}$ of $V^{\prime}$ verifies whether in all copies there is a reason the simulated maximal runs do not accept.
Recall that for a run there are three scenarios in which it is not accepted: (1) it reaches the end of the word, but the reached configuration is not accepted, (2) at some moment it tries to decrease some counter below zero, and (3) at some moment there is no transition available over the input letter. In the proof of Theorem 15 it was shown how a VASS can handle all three reasons. In short words: in case (1) it simulates the run till the end of the word and then checks that the reached configuration is not accepting and in cases (2) and (3) it guesses the moment in which there is no valid transition available and keeps this configuration untouched (in other words freezes it) till the end of the run when it checks by the accepting condition that the guess was correct. We only sketch how the downward-VASS $V^{\prime}$ works without stating explicitly its states and transitions. It starts in the configuration $c^{\prime}_{I}$ which consists of $k$ copies of $c_{I}$. Then it simulates the run in all the copies in the same way until the first time when there is a choice of transition. Then we enforce that at least one copy follows each choice, but we allow for more than one copy to follow the same choice. In the state of $V^{\prime}$ we keep the information which copies follow the same maximal run and which have already split. Each copy is exactly as in the proof of Theorem 15, it realises one of the scenarios (1), (2) or (3). As we know that $V$ is $k$-deterministic, we are sure that all the possible runs of $V$ can be simulated by $V^{\prime}$ under the condition that $V^{\prime}$ correctly guesses which copies should simulate which runs. If the guesses of $V^{\prime}$ are wrong and at some point it cannot send to each branch a copy, then the run of $V^{\prime}$ is rejected.
@@
The inclusion problem of a VASS language in a $k$-deterministic VASS language is in Ackermann.
@@ 4. Unambiguous VASS
Theorem 24 in [CLM+18] says that there exists a regular separator of $L_{1}$ and $L_{2}$ of size at most triply-exponential. In order to compute it, we can simply enumerate all the possible separators of at most triply-exponential size and check them one by one. For a given regular language and a given VASS language by Proposition 11 one can check in doubly-exponential time whether they nonempty intersect.
@@
For any VASS one can compute in an elementary time a finite family of regular languages that separates its languages.
@@ 5. Boundedly-ambiguous VASS
In this section we aim to prove Theorem 2 and show that Theorem 3 is its easy consequence. Recall the statements of both Theorem 2 and Theorem 3, the formulation of the statement of Theorem 2 is slightly simplified here.
Theorem 2. For each $k\in\mathbb{N}$ and a $k$-ambiguous VASS one can build in elementary time a downward-VASS which recognises the complement of its language.
@@
For each $k$-ambiguous VASS $V$ one can construct in doubly-exponential time a language equivalent VASS $V^{\prime}$ with the property that its control automaton is $k$-ambiguous.
Now our aim is to obtain a $k$-deterministic VASS $V^{2}$, which is language equivalent to $V^{1}$. We are not able to achieve it literally, but using the notion of $(M,h)$-decoration from Section 4 we can compute a somehow connected $k$-deterministic VASS $V^{2}$. We use the following lemma which is proved in Section 5.2.
@@
Now we consider the control automaton $\mathcal{A}$ of VASS $V^{1}$. We compute a monoid $M$ together with a homomorphism $h:\Sigma^{*}\to M$ that recognises all the state languages of $\mathcal{A}$. Then we construct the automaton $\mathcal{A}_{(M,h)}$. Note that the decoration of a VASS produces an HVASS, but as we decorate an automaton i.e. 00-VASS we get a 00-HVASS which is also a finite automaton. Based on $\mathcal{A}_{(M,h)}$ we construct a VASS $V^{2}$. We add a vector to every transition in $\mathcal{A}_{(M,h)}$ to produce a VASS that recognises the $(M,h)$-decoration of the language of VASS $V^{1}$. Precisely, if we have a transition $((p,m),(a,m^{\prime}),(q,m^{\prime}))$ in $\mathcal{A}_{(M,h)}$ then it is created from the transition $(p,a,q)$ in $\mathcal{A}$, which originates from the transition $(p,a,v,q)$ in $V^{1}$. So in $V^{2}$ we label $((p,m),(a,m^{\prime}),(q,m^{\prime}))$ with $v$ i.e. we have the transition $((p,m),(a,m^{\prime}),v,(q,m^{\prime}))$. Similarly, based on $V^{1}$, we define initial and final configurations in $V^{2}$. It is easy to see that there is a bijection between accepting runs in $V^{1}$ and accepting runs in $V^{2}$. By Lemma 30 $\mathcal{A}_{(M,h)}$ is $k$-deterministic which immediately implies that $V^{2}$ is $k$-deterministic as well.
@@ 5.1. Proof of Lemma 29
Lemma 29. For each $k$-ambiguous VASS $V$ one can construct in doubly-exponential time a language equivalent VASS $V^{\prime}$ with the property that its control automaton is $k$-ambiguous.
@@
Let $V=(\Sigma,Q,T,c_{I},F)$ be a $k$-ambiguous $d$-VASS for some $k\in\mathbb{N}$. We aim at constructing a language equivalent VASS $V^{\prime}$ such that its control automaton is $k$-ambiguous. Notice that if the control automaton of $V^{\prime}$ is $k$-ambiguous then clearly $V^{\prime}$ is $k$-ambiguous as well. The idea behind the construction of $V^{\prime}$ is that it behaves as $V$, but the additionally states of $V^{\prime}$ keep some finite information about the values of the counters. More precisely counter values are kept exactly until they pass some threshold $M$. After passing this threshold, its value is not kept in the state, only the information about passing the threshold is remembered. We define the described notion below more precisely as the $M$-abstraction of a VASS. The control state of $V^{\prime}$ is the $M$-abstraction of $V$ for appropriately chosen $M$; we show how to choose $M$ later.
@@
The proof of Lemma 31 follows the lines of the Rackoff construction for the ExpSpace algorithm deciding the coverability problem in VASS [Rac78]. One could tend to think that Lemma 31 can be proven just by application of [Rac78]. However, it seems that some adaptation of these techniques is necessary.
@@
The proof essentially follows the idea of the Rackoff construction [Rac78] showing that if there is a covering run in VASS then there is one of at most doubly-exponential length. It seems that it is not straightforward to apply the Rackoff construction as a black box, we need to repeat the whole argument in our setting. The reason is that we need to show that if there is a run in the $M$-abstraction of $V$, then there is also a run in $V$. However, some runs in the $M$-abstraction might correspond to runs of $V$, which are not valid, namely the counter values drop below zero. In order to show that we can modify such runs to obtain a valid runs of $V$ we adapt the Rackoff techniques, but we are not able to come up with a direct application of the Rackoff’s result.
Let $J\subseteq[1,d]$. We first define a $J$-restriction of $d$-dimensional vectors, configurations, and VASS. The $J$-restriction of a vector $v\in\mathbb{N}^{d}$ is a vector $v_{J}\in\mathbb{N}^{|J|}$ such that all the entries in coordinates not in $J$ are omitted. We generalise the notion of the $J$-restriction to configurations in a natural way: the $J$-restriction of a configuration $q(v)$ is the configuration $q(v_{J})$ where $v_{J}$ is the $J$-restriction of $v$. Now we define the $J$-restriction $V_{J}$ of VASS $V$. Let $V=(Q,T,c^{I},F)$ be a $d$-VASS, we omit the alphabet component, as it is not relevant here. Then $V_{J}=(Q,T_{J},c^{I}_{J},F_{J})$ is a $|J|$-dimensional VASS where
  • $(p,v_{J},q)\in T_{J}$ if there is $(p,v,q)\in T$ such that $v_{J}$ is the $J$-restriction of $v$,
  • $c^{I}_{J}$ is the $J$-restriction of $c^{I}$,
  • $c_{J}\in F_{J}$ if there is $c\in F$ such that $c_{J}$ is the $J$-restriction of $c$.
For a VASS $V$ let $\textup{norm}(V)$ be the maximal absolute value of the numbers occurring on transitions of $V$, in the initial configuration $c^{I}$ and in the representation of the set of final configurations $F$. Let us define a function $F$ taking two arguments: a $d$-VASS $V$ and a number $k\in[0,d]$ as follows $F(V,k)=(4|Q|\cdot(\textup{norm}(V)+1))^{(4d)^{k}}$. A key ingredient of the proof of Lemma 31 is the following proposition.
Proposition 32.
For any $d$-VASS $V$, subset of coordinates $J\subseteq[1,d]$ of size $k$, a configuration $c$ of $V_{J}$ and a number $M\in\mathbb{N}$ if there is an accepting run in the $M$-abstraction of $V_{J}$ starting from $c$ then there is also such a run of length bounded by $F(V,k)$.
Notice now that applying Proposition 32 to a $d$-VASS $V$, set $J=[1,d]$, the initial configuration $c^{I}$ of $V$ and a number $M=\textup{norm}(V)\cdot(F(V,d)+1)+1$ we get that if there is an accepting run in the $M$-abstraction of $V$ then there is also such a run $\rho$ of length bounded by $F(V,d)$. Maximal counter value on $\rho$ is bounded by maximal value in $c^{I}$ plus $\textup{norm}(V)\cdot F(V,d)$, namely by $\textup{norm}(V)\cdot(F(V,d)+1)$. This number is smaller than $M$ therefore $\rho$ is also a run in $V$, which finishes the proof of Lemma 31 as $F(V,d)$ is doubly-exponential.
Thus, it remains to show Proposition 32.
Proof 5.4 (Proof of Proposition 32).
We prove Proposition 32 by induction on the size of $J$. If $J=\emptyset$ then clearly the length of the shortest accepting run is bounded by number of states in $Q$, so also by $|Q|\leq 4|Q|\cdot(\textup{norm}(V)+1))=F(V,0)$.
For an induction step assume that Proposition 32 holds for $|J|\leq k$. We aim to prove it for $|J|=k+1$. Let us consider some $M$ and some configuration $c_{1}$. Suppose that
$\rho=(c_{1},t_{1},c^{\prime}_{1}),(c_{2},t_{2},c^{\prime}_{2}),\ldots,(c_{n},t_{n},c^{\prime}_{n})$
is the shortest accepting run starting from $c$ in the $M$-abstraction of $V_{J}$. Let $G(V,k)=\textup{norm}(V)\cdot(F(V,k)+1)$. There are two possibilities
  1. (1)
    for every $c_{i}$ on $\rho$ and every $j\in J$ holds $c_{i}[j]<G(V,k)$,
  2. (2)
    there exists $c_{i}$ on $\rho$ and $j\in J$ such that $c_{i}[j]\geq G(V,k)$ and $c_{i}$ is the first such configuration on $\rho$.
In the first case, the length of the run is smaller than $|Q|\cdot(G(V,k))^{k+1}$ as there are exactly $|Q|\cdot(G(V,k))^{k+1}$ configurations with counters bounded by $G(V,k)$. Thus, it remains to show that $|Q|\cdot(G(V,k))^{k+1}\leq F(V,k+1)$, we show it after considering the second case.
In the second case notice that from $c_{i}$ there is an accepting run in the $M$-abstraction of $V_{J}$. Thus from the $(J\setminus\{j\})$-restriction of $c_{i}$ there is an accepting run in the $M$-abstraction of $V_{(J\setminus\{j\})}$. By induction assumption, a shortest accepting run from the $(J\setminus\{j\})$-restriction of $c_{i}$ in the $M$-abstraction of $V_{(J\setminus\{j\})}$ is of length at most $F(V,k)$. Notice, that as $c_{i}[j]\geq G(V,k)=\textup{norm}(V)\cdot(F(V,k)+1)$ the corresponding run of length bounded by $F(V,k)$ can be performed from $c_{i}$ also in $V_{J}$. Moreover, this is an accepting run in $V_{J}$ as (1) on the $j$-th counter the run ends with the counter value at least $G(V,k)-\textup{norm}(V)\cdot F(V,k)\geq\textup{norm}(V)$ which is bigger than any constant in the description of the set $F_{J}$, and (2) on other counters we use the fact the the run was accepting in $V_{J\setminus\{j\}}$.
Thus in the second case the length of the run can be bounded by the length of the part before $c_{i}$ plus the length of the part after $c_{i}$ which is bounded by $F(V,k)$. The length of the run before $c_{i}$ is bounded by $|Q|\cdot(G(V,k))^{k+1}$ similarly as in the first case. Thus it remains to show that
$F(V,k)+|Q|\cdot(G(V,k))^{k+1}\leq F(V,k+1).$ (1)
This inequality implies the inequality which remained to be shown in the first case, so we finish the proof by showing estimation (1). Let us denote $K=\textup{norm}(V)$. Recall that $F(V,k)=(4|Q|\cdot K)^{(4d)^{k}}$ and $G(V,k)=K\cdot(F(V,k)+1)$. Therefore
$F(V,k)+|Q|\cdot(G(V,k))^{k+1}$ $\leq 2|Q|\cdot(G(V,k))^{k+1}$
$\leq 2|Q|\cdot(2K\cdot F(V,k))^{k+1}$
$\leq(4|Q|\cdot K\cdot F(V,k))^{k+1}$
$\leq(4|Q|\cdot K\cdot(4|Q|\cdot K)^{(4d)^{k}})^{k+1}$
$\leq((4|Q|\cdot K)^{((4d)^{k}+1)})^{k+1}$
$\leq(4|Q|\cdot K)^{((4d)^{k}+1)\cdot(k+1)}$
$\leq(4|Q|\cdot K)^{4d\cdot(4d)^{k}}$
$\leq(4|Q|\cdot K)^{(4d)^{k+1}}=F(V,k+1),$
which finishes the proof.

### v4

@@ 3.2. Upper bounds
There is an algorithm that, for a $k$-control-deterministic $d$-VASS $V$ constructs, in exponential time, a $(k\cdot d)$-dimensional downward-VASS which recognises the complement of the language of $V$.
@@
Before starting the proof, let us remark that it would seem natural to first build a deterministic $(k\cdot d)$-VASS equivalent to the input $k$-control-deterministic $d$-VASS and then apply construction from the proof of Theorem 15 to recognise its complement. However, it is not clear how to construct a deterministic $(k\cdot d)$-VASS equivalent to $k$-control-deterministic $d$-VASS, thus we compute directly a VASS recognising the complement of the input VASS language.
Let $V=(\Sigma,Q,T,c_{I},F)$ be a $k$-control-deterministic $d$-VASS. We aim to construct a $(k\cdot d)$-dimensional downward-VASS $V^{\prime}=(\Sigma,Q^{\prime},T^{\prime},c^{\prime}_{I},F^{\prime})$ such that $L(V^{\prime})=\Sigma^{*}\setminus L(V)$. In this proof, also, we strongly rely on the ideas introduced in the proof of Theorem 15. Recall that the idea of the construction is that $V^{\prime}$ simulates $k$ copies of $V$ which take care of different maximal runs of $V$. Then the accepting condition $F^{\prime}$ of $V^{\prime}$ verifies whether in all copies there is a reason the simulated maximal runs do not accept.
Recall that for a run there are three scenarios in which it is not accepted: (1) it reaches the end of the word, but the reached configuration is not accepted, (2) at some moment it tries to decrease some counter below zero, and (3) at some moment there is no transition available over the input letter. In the proof of Theorem 15 it was shown how a VASS can handle all three reasons. In short words: in case (1) it simulates the run till the end of the word and then checks that the reached configuration is not accepting and in cases (2) and (3) it guesses the moment in which there is no valid transition available and keeps this configuration untouched (in other words freezes it) till the end of the run when it checks by the accepting condition that the guess was correct. We only sketch how the downward-VASS $V^{\prime}$ works without stating explicitly its states and transitions. It starts in the configuration $c^{\prime}_{I}$ which consists of $k$ copies of $c_{I}$. Then it simulates the run in all the copies in the same way until the first time when there is a choice of transition. Then we enforce that at least one copy follows each choice, but we allow for more than one copy to follow the same choice. In the state of $V^{\prime}$ we keep the information which copies follow the same maximal run and which have already split. Each copy is exactly as in the proof of Theorem 15, it realises one of the scenarios (1), (2) or (3). As we know that $V$ is $k$-control-deterministic, we are sure that all the possible runs of $V$ can be simulated by $V^{\prime}$ under the condition that $V^{\prime}$ correctly guesses which copies should simulate which runs. If the guesses of $V^{\prime}$ are wrong and at some point it cannot send to each branch a copy, then the run of $V^{\prime}$ is rejected.
@@
The inclusion problem of a VASS language in a $k$-control-deterministic VASS language is in Ackermann.
@@ 4. Unambiguous VASS
Theorem 24 in [CLM+18] says that there exists a regular separator of $L_{1}$ and $L_{2}$ of size at most triply-exponential. In order to compute it, we can simply enumerate all the possible separators of at most triply-exponential size and check them one by one. For a given regular language and a given VASS language by Proposition 11 one can check in doubly-exponential time whether they intersection is nonempty.
@@
There is an algorithm that, for any VASS computes, in an elementary time, a finite family of regular languages that separates its languages.
@@ 5. Boundedly-ambiguous VASS
In this section we aim to prove Theorem 2 and show that Theorem 3 is its easy consequence. Recall the statements of both Theorem 2 and Theorem 3.
Theorem 2. There is an algorithm that, for each $k\in\mathbb{N}$ and a $k$-ambiguous VASS $V$ constructs, in elementary time, a VASS with a downward-closed set of accepting configurations which recognises the complement of the language of $V$.
@@
There is an algorithm, that for each $k$-ambiguous VASS $V$ constructs, in doubly-exponential time, a language equivalent VASS $V^{\prime}$ with the property that its control automaton is $k$-ambiguous.
Now our aim is to obtain a $k$-control-deterministic VASS $V^{2}$, which is language equivalent to $V^{1}$. We are not able to achieve it literally, but using the notion of $(M,h)$-decoration from Section 4 we can compute a somehow connected $k$-control-deterministic VASS $V^{2}$. We use the following lemma which is proved in Section 5.2.
@@
Now we consider the control automaton $\mathcal{A}$ of VASS $V^{1}$. We compute a monoid $M$ together with a homomorphism $h:\Sigma^{*}\to M$ that recognises all the state languages of $\mathcal{A}$. Then we construct the automaton $\mathcal{A}_{(M,h)}$. Note that the decoration of a VASS produces an HVASS, but as we decorate an automaton i.e. 00-VASS we get a 00-HVASS which is also a finite automaton. Based on $\mathcal{A}_{(M,h)}$ we construct a VASS $V^{2}$. We add a vector to every transition in $\mathcal{A}_{(M,h)}$ to produce a VASS that recognises the $(M,h)$-decoration of the language of VASS $V^{1}$. Precisely, if we have a transition $((p,m),(a,m^{\prime}),(q,m^{\prime}))$ in $\mathcal{A}_{(M,h)}$ then it is created from the transition $(p,a,q)$ in $\mathcal{A}$, which originates from the transition $(p,a,v,q)$ in $V^{1}$. So in $V^{2}$ we label $((p,m),(a,m^{\prime}),(q,m^{\prime}))$ with $v$ i.e. we have the transition $((p,m),(a,m^{\prime}),v,(q,m^{\prime}))$. Similarly, based on $V^{1}$, we define initial and final configurations in $V^{2}$. It is easy to see that there is a bijection between accepting runs in $V^{1}$ and accepting runs in $V^{2}$. By Lemma 30 $\mathcal{A}_{(M,h)}$ is $k$-deterministic which immediately implies that $V^{2}$ is $k$-control-deterministic as well.
@@ 5.1. Proof of Lemma 29
Lemma 29. There is an algorithm, that for each $k$-ambiguous VASS $V$ constructs, in doubly-exponential time, a language equivalent VASS $V^{\prime}$ with the property that its control automaton is $k$-ambiguous.
@@
Let $V=(\Sigma,Q,T,c_{I},F)$ be a $k$-ambiguous $d$-VASS for some $k\in\mathbb{N}$. We aim at constructing a language equivalent VASS $V^{\prime}$ such that its control automaton is $k$-ambiguous. Notice that if the control automaton of $V^{\prime}$ is $k$-ambiguous then clearly $V^{\prime}$ is $k$-ambiguous as well. The idea behind the construction of $V^{\prime}$ is that it behaves as $V$, but additionally the states of $V^{\prime}$ keep some finite information about the values of the counters. More precisely counter values are kept exactly until they pass some threshold $M$. After passing this threshold, its value is not kept in the state, only the information about passing the threshold is remembered. We define the described notion below more precisely as the $M$-abstraction of a VASS. The control state of $V^{\prime}$ is the $M$-abstraction of $V$ for appropriately chosen $M$; we show how to choose $M$ later.
@@
Lemma 31 is a consequence of the result by Rackoff [Rac78] showing that if there is a covering path in VASS then there is also a covering path of at most doubly-exponential length.
@@
Let $V=(\Sigma,Q,T,c_{I},F)$ and let ${\mathcal{C}}$ be the set of configurations of $V$ from which it is possible to reach $F$. As the set $F$ is upward closed the same holds for ${\mathcal{C}}$. Further, let $\bar{{\mathcal{C}}}$ be $(\mathbb{N}\cup\omega)^{d}\setminus{\mathcal{C}}$, so the complement of ${\mathcal{C}}$ in the space of $(\mathbb{N}\cup\omega)^{d}$. It is downward closed i.e. it is a union of down-atoms, $d_{1}\downarrow,d_{2}\downarrow,\ldots,d_{n}\downarrow$. Let $||V||$ be the norm of VASS $V$, namely the maximal absolute value occurring on transition of $V$. Let $M=\max\{d_{i}[j]:i\leq n,j\leq d,d_{i}[j]\neq\omega\}+||V||$ be the maximal constant that appears in the description of the set $\bar{{\mathcal{C}}}$ plus the norm of VASS $V$. By [Rac78] the minimal length of a path covering a given configuration is at most doubly-exponential wrt. the size of the VASS and the norm of the target configuration (i.e. the maximal absolute value of coordinates in the target). Therefore all the numbers in vectors $d_{i}$ are at most doubly-exponential as well. Thus also the constant $M$ is at most doubly-exponential.
By an $M$-abstraction of a configuration $c$, denoted by $A_{M}(c)$, we mean an $\omega$-configuration in $Q\times(\mathbb{N}\cup\{\omega\})^{d}$ where we substitute in $c$ any value greater or equal $M$ by $\omega$. Concretely:
$A_{M}(c)[j]=\begin{cases}c[j]&\text{if }c[j]<M,\\ \omega&\text{if }c[j]\geq M.\end{cases}$
Observe that function $A_{M}$ has a following property, which is the consequence of the definition of the value $M$,
$c\in\bar{{\mathcal{C}}}\iff A_{M}(c)\in\bar{{\mathcal{C}}}.$ (1)
Finally, the image of the function $A_{M}$ is in one to one correspondence with the states of the $M$-abstraction of $V$, so we will use $M$-abstractions of configurations in both settings: in the VASS $V$ and in the $M$-abstraction of $V$.
Now we can start the actual proof of the lemma. We prove it by contradiction. Namely we assume that from some configuration $c$ there is no accepting run in $V$, but there is an accepting run from the state $A_{M}(c)$ in the $M$-abstraction of $V$. Therefore $c\in\bar{{\mathcal{C}}}$ and from $A_{M}(c)$ in the $M$-abstraction it is possible to reach the set of final states $F^{\prime}$ of the $M$-abstraction. We will show that such $c$ cannot exist. First observe that $A_{M}(c)$ cannot be among final states of the $M$-abstraction as then $c\in F$ which contradicts the fact that $c\in\bar{{\mathcal{C}}}$.
So for every $A_{M}(c)$ there is a shortest path in the $M$-abstraction to the set of final states $F^{\prime}$. Without loss of generality we can assume that $c$ is chosen in such a way that the length of the path from $A_{M}(c)$ to $F^{\prime}$ is minimal. We denote this path by $\sigma$. As $A_{M}(c)\not\in F^{\prime}$ we can decompose $\sigma=t\sigma^{\prime}$ into the head and a tail, where $t=(q,a_{t},\delta_{t},p)$. Let us consider $\omega$-configuration $c^{\prime}=A_{M}(c)+\textup{eff}(t)$ denoting an $\omega$-configuration $(p,v)$ where $v[i]=A_{M}(c)[i]+\delta_{t}[i]$ for every $i\in[1,d]$ On the one hand we know that $c^{\prime}\in\bar{{\mathcal{C}}}$, as $c\in\bar{{\mathcal{C}}}$. Because of (1) we get that $A_{M}(c^{\prime})\in\bar{{\mathcal{C}}}$. On the other hand in the $M$-abstraction of $V$ the path $\sigma$ from $A_{M}(c)$ traverses through $A_{M}(c^{\prime})$. Therefore the path $\sigma^{\prime}$ from $A_{M}(c^{\prime})$ to $F^{\prime}$ is shorter than the path $\sigma$. So we get a contradiction with the choice of $c$ as $c^{\prime}\in\bar{{\mathcal{C}}},$ and the path from $A_{M}(c^{\prime})$ to $F^{\prime}$ is shorter than from $c$.